OpenStudy (anonymous):

can someone please tell me how to solve this differential equation. y^6 + y = 0 . :(

OpenStudy (anonymous):

is that the differentiated equation or do you need to differentiate it?

OpenStudy (anonymous):

its a question of differential equations. i need to find the roots and then the solution of this equation

OpenStudy (anonymous):

okay differentiate it. you will get 6y^5 + 1 = 0 y^5 = -1/6 y = fifth roots of (-1/6)